A. Find the length of the following confidence interval.




1. Upper limit = 0.995




Lower Limit = 0.437



2. Upper limit = 394.14




Lower Limit = 354.74



3. Upper limit = 0.02946




Lower Limit = 0.02244



4. 0.475 < p < 0.735



5. 0.355 < p < 0.570

1
Expert's answer
2022-05-09T16:09:31-0400

1.


Lenght=0.9950.437=0.558Lenght=0.995-0.437=0.558

2.


Lenght=394.14354.74=39.40Lenght=394.14-354.74=39.40

3.


Length=0.029460.02244=0.00702Length=0.02946-0.02244=0.00702

4.


Lenght=0.7350.475=0.260Lenght=0.735-0.475=0.260

5.


Lenght=0.5700.355=0.215Lenght=0.570-0.355=0.215